OpenStudy (anonymous):

Multiply or divide as indicated. r^ 2* r *r^ 5

OpenStudy (anonymous):

Is this what the question looks like? \[r^2 * r * r^5\]

OpenStudy (anonymous):

* being multiplication

OpenStudy (anonymous):

yeah.

OpenStudy (anonymous):

\[r \times r \times r \times r \times r \times r \times r \times r\]

OpenStudy (anonymous):

r^8

OpenStudy (anonymous):

yep